# Edge AI: A Taxonomy, Systematic Review and Future Directions

### v1

@@ II-E Smart Healthcare
Intelligent healthcare systems focus on integrating AI and edge computing and the challenges related to privacy and security, decision-making, and optimization. These systems make use of technologies, including genetic-based encryption for data security, federated learning in the Internet of Medical Things, and nanosensor-equipped systems to improve efficiency and security. Mobile computing has played a vital role in healthcare, particularly during the COVID-19 pandemic, enabling telemedicine and contact tracing, emphasizing the significance of technology in tackling healthcare challenges. The following research papers delve deeper into the intersection of technology, healthcare, and privacy. In 2018, the authors [45] proposed an intelligent home monitoring system based on edge-fog computing with AI capabilities. Latency issues and reliability are the main concerns for the authors in developing smart home real-time applications. Further, the authors [46] provide a comprehensive overview of the key elements of the MCPS from multiple perspectives, covering design, methodology, and significant supporting technologies such as cloud computing, edge computing, IoT, sensor networks, and systems with multiple agents. In [47], authors provided a distinctive and specialized route resource recommendation (R3) protocol to handle resource management and connection problems in autonomous, connected ambulances (ACA) for route optimization. Furthermore, the authors [48] provide a condition-aware analytical framework that may be used to recommend health conditions in the IoT-based mobile healthcare systems. This framework corresponds to IoT devices that have limited resources, such as those with a memory utilization rate of 6.6%. Moreover, the authors [11] Present a flexible working mechanism that allows the combined configuration of data quality ratios (DQRs) and model complexity ratios (MCRs) for AI tasks and addresses a flexible model deployment architecture for edge AIaaS. In [49], researchers suggest using an edge-of-things (EoT) framework to implement centralized and federated transfer learning (CMTL) for cyberattack detection systems in the healthcare industry. Finally, the authors [50] presents a new approach called CoDoC, which stands for complementary-driven deferral-to-clinical workflow. Its purpose is to decide when to rely on a diagnostic AI model and when to hand it off to a clinician.
TABLE I: Comparison of Our Survey with Existing Related Surveys
Work Research Domain [2] [4] [6] [7] [8] [9] [10] [11] [12] [14] [16] Proposed work
@@
II-F Comparison with Related Surveys
@@
Table I shows the comparison of our systematic review with related surveys by focusing on the advancement of AI at the edge, optimizing algorithms in constrained environments, solving training data scarcity with AI techniques, and using AI/ML for resource management in fog and edge computing settings. There, the different columns are set on the basis of various applications, their methods of optimization, and fruitful utilization. This survey covers several edge computing concepts and emphasises the application of AI on edge devices with constrained resources. The survey examines the optimisation of ML algorithms for such restricted environments and covers current IoT applications across several sectors, including industrial automation, smart homes, and autonomous vehicles. It also highlights the difficulties and possible paths for edge computing and edge AI research, offering a strong basis for further investigation in the field. Considering the dynamic, diversified, and resource-constrained character of fog and edge computing settings, this research emphasises the potential of AI/ML, particularly reinforcement learning techniques, in addressing resource management challenges in these systems [2] [4] [6] [7]. A thorough analysis of the literature was done to look at how AI/ML applications may be used to effectively manage resources in these kinds of situations. A taxonomy was established to categorise and contrast different approaches. Enhancing explainability, reducing variance, and boosting online training of AI/ML algorithms are highlighted as critical future research directions to adapt to the constantly changing fog/edge computing landscapes. The study emphasises the significance and changing challenges of resource management [8], whereas a presentation framework is described that addresses the issue of sparse training data in emerging radio access networks by utilizing a range of methods, such as interpolation, domain-knowledge techniques, generative adversarial networks (GANs), transfer learning, autoencoders, few-shot learning, simulators, and testbeds. The challenges are highlighted and presented by insufficient training data, and the crucial role that Automation powered by AI plays in the operation, optimisation, and troubleshooting of cellular networks is described. The technique suggests an integrated strategy to improve data availability in cellular networks and includes a survey and taxonomy of current approaches to lessen this scarcity. In addition, the study emphasises the necessity for scalable, reliable solutions that take conditional contexts into account for generating high-dimensional data in radio access network applications also [9] [10]. Table I reveals a dearth of reviews on Edge AI, with most existing surveys and review papers presenting an overview or vision of the technology rather than a comprehensive survey, systematic review, and detailed taxonomy. To the best of the author’s knowledge, this is the first survey paper on edge AI that provides a thorough taxonomy and a systematic review and highlights future research topics.
III Edge AI: Background and Current Status
This section explains some concepts related to background and current status in Edge AI for the reader. Subsection III-A explains edge computing and its historical emergence. Subsection III-B provides information on the integration of AI and edge technologies. This section is completed by explaining Edge AI applications and challenges in subsection III-C and subsection III-D, respectively.
III-A Historical Emergence of Edge Computing
The concept of edge computing is a paradigm that brings computing resources closer to the data source, unlike the cloud, which provides services through a remote server [51]. In this way, it is aimed to reduce problems such as unnecessary bandwidth occupation and latency in today’s world where huge amounts of data that need to be processed are produced [52]. To understand the emergence of edge computing, it will be more useful to examine previous paradigms such as cloud and fog computing. Figure 1 shows the advantages of cloud, fog, and edge computing over each other and their layer arrangement. Now let’s briefly touch on these concepts:
Figure 1: Computing Paradigms and their Objectives.
  • Cloud Computing: Cloud computing is a paradigm that dates back to the 1970s and refers to the use of common computing resources by users on a server via the Internet [53]. Today, it is offered to users with various service models, especially by large companies such as Microsoft Azure, Google Cloud Platform and IBM Cloud. The advantages of cloud computing are as follows [54]:
    • Scalability, in case the need for computing resources increases (demand fluctuations), cloud computing provides services such as more processing power and storage by scaling the resources. In this way, performance measures such as SLA and QoS are ensured.
  • Fog Computing: The concept of fog computing was introduced by Cisco in 2012 [55]. This paradigm recommends moving computing resources closer to the endpoints of the network (such as routers and gateways) to reduce the latency and bandwidth problems that occur in cloud computing. When Figure 1 is examined, fog computing acts as a layer between the cloud and the edge. The advantages of fog computing are as follows [56]:
III-B Integration of AI with Edge Technology
The idea of combining the concepts of AI and edge computing is one of the hottest topics of recent times. Fig. 2 shows a main diagram explaining the concept of Edge AI. Edge AI is processing data in AI algorithms distributed on edge nodes (IoT and mobile devices) close to the data source, rather than processing data on a central cloud platform [60]. The advantages offered by these two technologies can be listed as follows [6]:
  • The amount of latency is critical in applications such as low latency, time-sensitive autonomous vehicles, and e-health where patients are monitored. With Edge and AI integration, this problem can be overcome by processing data in real time.
  • In the data security, edge paradigm, since the data will be processed at a point close to the source, sensitive data such as biometric data is processed on local devices within the network. In this way, security possibilities such as data leakage can be eliminated.
  • Scalable resources, edge paradigm consists of nodes consisting of homogeneous or heterogeneous devices. In this way, task scheduling can be made according to processing power needs.
Figure 2: Architecture of Edge AI.
III-C Edge AI Applications
Edge AI applications, created by combining the concepts of Edge and AI, provide lower latency and higher security than Cloud-based AI applications. Figure 3 shows some of these applications. Let’s explain these applications:
  • Smart Parking: With the increase in means of transportation, parking has become a big problem, especially in big cities. Edge AI-based solutions with the help of sensors and IoTs can be used to solve these problems [61].
Figure 3: The Edge AI Applications.
Figure 4: Edge AI Implementation Challenges.
III-D Edge AI Implementation Challenges
EdgeAI, which emerges by combining Edge and AI, brings with it the advantages it offers, but also challenges that are still waiting to be solved. These challenges are shown in Figure 4. Let’s explain these challenges:
  • Energy Efficiency: Edge devices generally consist of homogeneous and heterogeneous devices with low processing and storage capacity. Applications that require Natural Language Processing (NLP) and intensive image processing will cause excessive resource consumption on edge devices [64]. For this reason, new solutions such as special AI chips or task engineering are needed.
  • Maintenance and Updates: Since edge devices consist of devices distributed in different locations, this means more attack targets for hackers [65]. In addition, not all devices in the edge nodes have a homogeneous structure, which means separate system maintenance and updates for each node [66]. Measures such as automatic updating can be taken to solve these problems.
  • Scalability: Since edge devices generally consist of heterogeneous devices, the distribution of a single application to different devices is still a challenge (task scheduling, etc.) [67]. Additionally, it is difficult to synchronize data across all devices. Effective microservice architectures and load-balancing algorithms that prevent a node from being overloaded can be used to solve this problem.
@@ IV Review Methodology
This article does a systematic review to categorise studies that are pertinent to this study domain or discuss specific research questions on ”Edge AI”. In this article, we used the guidelines established by Kitchenham et al. [68, 69, 70] to provide a comprehensive review of Edge AI. The review is the optimal and reliable approach to document and analyse current research works. The systematic approach enables researchers to carefully analyse the positive and negative aspects of recent studies, conduct a thorough examination to identify potential gaps in research and future trends and difficulties, and provide a solid foundation and context for establishing a new study field. Furthermore, the complete research approach is presented in Figure 5, that represents the structure of the process that was used in the systematic review research.
@@ IV-A Design and Plan of Review
The review procedure illustrates the methodologies applied to conduct a systematic review with the aim of minimizing the potential for biased research. Therefore, possessing a pre-established process is crucial. In the absence of a systematic methodology, researchers’ predispositions have the ability to influence the process of selecting and analyzing studies. This may result in the omission of crucial inquiries essential for a thorough analysis and comprehension of the subject matter. The review process encompasses the research inquiries, exploration approach, criteria for selecting studies, procedures for assessing quality, and techniques for extracting and synthesizing data [71].
@@ IV-B Research Questions
Determining the research queries is crucial in the method of planning to develop a strong systematic review. The design of the research problem requires a thorough examination of existing literature studies. The primary aim of the present systematic review is to thoroughly examine and evaluate the various methods and strategies being employed for the  edge intelligence or edge AI. Furthermore, in order to emphasise the research findings and effectively showcase the useful consequences, the research questions that follow have been defined in Table II.
@@
RQ1 What are the primary methodologies and strategies adopted in Edge AI, and how do they vary in terms of applications, infrastructure, and algorithms? The purpose of this research question is to understand the various methods and approaches used in Edge AI. Infrastructure, Application Architecture, and Methods Sections 5.1, 5.2, and 5.4
RQ2 To what extent may heuristic algorithms enhance the effectiveness and performance of Edge AI systems, specifically in tasks such as model training and optimization? This research question helps in examining the role of heuristic methods in Edge AI. Heuristic Methods Section 5.4.1
RQ3 How do the meta-heuristic algorithms used by Edge AI solve difficult optimization problems, and what are their applications and prime characteristics? The characteristics and applications of meta-heuristics methods are investigated in this paper. Meta-Heuristics Methods Section 5.4.2
RQ4 How can real-time data processing and decision-making in Edge AI devices get aided by ML techniques, including deep learning models? This research question analyzes the role of ML techniques. Machine Learning Methods Section 5.4.3
RQ5 What are the main benefits and constraints of Edge AI’s deep reinforcement learning, specifically for situations like autonomous driving? The pros and cons of deep reinforcement learning are examined via this research question. Deep Reinforcement Learning Methods Section 5.4.4
RQ6 How do Edge AI devices deal with heterogeneity in computational, hardware, and platform factors, and what approaches are used to optimize resource allocation? Heterogeneity in various aspects is discussed along with optimization of resource allocation approaches. Heterogeneity and Horizontal, Vertical, and Hybrid container scaling Sections 5.7 and 5.11.2
RQ7 What are the main challenges and measures for implementing Edge AI applications, specifically in terms of resource provisioning and workload distribution? The purpose of this research question was to identify challenges and solutions while implementing Edge AI applications. Resource Provisioning and Workload Distribution and Prediction Sections 5.5.1 and 5.5.4
RQ8 What are the pros and cons of both monolithic and microservices architectures in terms of Edge AI flexibility, performance, and resource utilization? This research question helps in comparing both monolithic and microservice architectures. Comparisons of Existing Edge AI approaches on the basis of Application Section 6.2
RQ9 What are the primary factors worth considering for resource management in Edge AI including resource provisioning, allocation, application placement, workload distribution, and prediction? The prime factors for resource management are examined through this research question. Resource Management Section 5.5
RQ10 How do ML model sizing impact the performance and resource consumption of Edge AI applications? This research question investigates the impact of machine learning model size on the performance and resource consumption of Edge AI applications. Full ML Model Sizing Section 5.6.2
RQ11 What are the key components that affect the choice of Edge AI infrastructure, such as Cloud, Fog, and Edge computing, with regard to how they affect application performance and resource utilization? This research question helps identify components that influence the choice of Edge AI infrastructure. Infrastructure Section 5.1
@@ IV-C2 Search Query
A comprehensive search was conducted utilising Logical OR/AND operators to connect the keywords, concepts, synonyms, and abbreviations. The initial phase entails conducting an automated search using predetermined keywords that align with the study topics of this systematic literature review (SLR). The keywords used are [(((”Edge AI” OR ”Edge Intelligence” OR ”Edge Computing”) AND (”Machine Learning” OR ”Deep Learning” OR ”Reinforcement Learning”)) AND (”Resource Management” OR “Resource Allocation”)) OR (“Cloud Computing” OR “Fog Computing”) OR (“Application Placement” OR “Security” OR “Scheduling” OR “Simulation”)]. The search terms are obtained from the specified research topics and the framework of this systematic review in order to encompass the most significant and interrelated publications.
@@ V A Taxonomy of Edge AI
Based on the current advancements and existing literature on edge AI, we have proposed a taxonomy, as shown in Fig. 7.
@@ V-A Infrastructure
AI models and applications can be deployed in different infrastructures based on their target scenarios. The three most common infrastructures widely used are Cloud, Fog and Edge.
@@ V-A1 Cloud
The essence of cloud computing is to pool resources with virtualization technology. Virtualization technology transforms a physical machine into several virtual machines, greatly changing the mode of application operation and deployment. The term ’cloud’ here refers to remote data centers [21]. Users usually connect to the servers in the data center through the network to use the computing resources here. The essence of cloud computing is to pool resources, allowing users to purchase resources according to their own requirements, greatly reducing resource waste.
@@ V-A2 Fog
As is well known, fog is closer to the ground than cloud, and the ground here refers to the user’s local device. Fog computing technology adopts distributed computing technology, which arranges computing resources on the side closer to the user than cloud computing [6]. It can be said that fog computing has broadened the network computing mode of cloud computing, widening computing capacity from the network center to the network edge, and thus more widely applied to numerous services. Fog computing is more widely distributed geographically and has greater mobility, making it suitable for an increasing number of intelligent devices that do not require extensive computation. For some time delay-sensitive applications such as real-time interaction system, fog computing also has greater advantages.
Image, video and natural language processing (NLP) are the recently emerging applications of fog computing. The placement and processing of images in fog computing is one of the most widely used fields of AI in research and industry. Its goal is to distinguish objects and people from each other and classify and distinguish photos based on image processing algorithms. Using fog computing in image processing-based applications can shorten response delay and improve service quality. In medical applications that require image processing accuracy and fast processing of medical data, deploying effective scheduling algorithms in foggy environments may be beneficial. According to other works, DL algorithms such as Generative Adversarial Networks (GAN) and Convolutional Neural Networks (CNN) can commonly be used in the field of image processing in fog.

### v2

@@ III-A5 Smart Healthcare
Intelligent healthcare systems focus on integrating AI and edge computing and the challenges related to privacy and security, decision-making, and optimization. These systems make use of technologies, including genetic-based encryption for data security, federated learning in the Internet of Medical Things, and nanosensor-equipped systems to improve efficiency and security. Mobile computing has played a vital role in healthcare, particularly during the COVID-19 pandemic, enabling telemedicine and contact tracing, emphasizing the significance of technology in tackling healthcare challenges. The following research papers delve deeper into the intersection of technology, healthcare, and privacy. In 2018, the authors [66] proposed an intelligent home monitoring system based on edge-fog computing with AI capabilities. Latency issues and reliability are the main concerns for the authors in developing smart home real-time applications. Further, the authors [67] provide a comprehensive overview of the key elements of the MCPS from multiple perspectives, covering design, methodology, and significant supporting technologies such as cloud computing, edge computing, IoT, sensor networks, and systems with multiple agents. In [68], authors provided a distinctive and specialized route resource recommendation (R3) protocol to handle resource management and connection problems in autonomous, connected ambulances (ACA) for route optimization. Furthermore, the authors [69] provide a condition-aware analytical framework that may be used to recommend health conditions in IoT-based mobile healthcare systems. This framework corresponds to IoT devices that have limited resources, such as those with a memory utilization rate of 6.6%. Moreover, the authors [11] Present a flexible working mechanism that allows the combined configuration of data quality ratios (DQRs) and model complexity ratios (MCRs) for AI tasks and addresses a flexible model deployment architecture for edge AIaaS. In [70], researchers suggest using an edge-of-things (EoT) framework to implement centralized and federated transfer learning (CMTL) for cyberattack detection systems in the healthcare industry. Finally, the authors [71] presents a new approach called CoDoC, which stands for complementary-driven deferral-to-clinical workflow. Its purpose is to decide when to rely on a diagnostic AI model and when to hand it off to a clinician.
TABLE I: Comparison of Our Survey with Related Surveys
Work Research Domain [2] [4] [6] [7] [8] [9] [10] [11] [12] [14] [16] Our Survey (this paper)
@@
III-B Related Surveys and Our Contributions
Table I shows the comparison of our systematic review with related surveys by focusing on the advancement of AI at the edge, optimizing algorithms in constrained environments, solving training data scarcity with AI techniques, and using AI/ML for resource management in fog and edge computing settings. There, the different columns are set on the basis of various applications, their methods of optimization, and fruitful utilization. This survey covers several edge computing concepts and emphasises the application of AI on edge devices with constrained resources. The survey examines the optimisation of ML algorithms for such restricted environments and covers current IoT applications across several sectors, including industrial automation, smart homes, and autonomous vehicles. It also highlights the difficulties and possible paths for edge computing and edge AI research, offering a strong basis for further investigation in the field.
III-B1 Related Surveys
@@
Considering the dynamic, diversified, and resource-constrained character of fog and edge computing settings [2], this research emphasises the potential of AI/ML, particularly reinforcement learning techniques [4], in addressing resource management challenges in these systems [6]. A thorough analysis of the literature was done to look at how AI/ML applications may be used to effectively manage resources in these kinds of situations [11]. A taxonomy was established to categorise and contrast different approaches [7]. Enhancing explainability, reducing variance, and boosting online training of AI/ML algorithms are highlighted as critical future research directions to adapt to the constantly changing fog/edge computing landscapes [16]. The study emphasises the significance and changing challenges of resource management [8], whereas a presentation framework is described that addresses the issue of sparse training data in emerging radio access networks by utilizing a range of methods, such as interpolation, domain-knowledge techniques, generative adversarial networks (GANs), transfer learning, autoencoders, few-shot learning, simulators, and testbeds [12]. The challenges are highlighted and presented by insufficient training data, and the crucial role that Automation powered by AI plays in the operation, optimisation, and troubleshooting of cellular networks is described [14]. The technique suggests an integrated strategy to improve data availability in cellular networks and includes a survey and taxonomy of current approaches to lessen this scarcity. In addition, the study emphasises the necessity for scalable, reliable solutions that take conditional contexts into account for generating high-dimensional data in radio access network applications also [9] [10]. Table I reveals a dearth of reviews on Edge AI, with most existing surveys and review papers [2] [4] [6] [7] [8] [9] [10] [11] [12] [14] [16] presenting an overview or vision of the technology rather than a comprehensive survey, systematic review, and detailed taxonomy. To the best of the author’s knowledge, this is the first survey paper on edge AI that provides a thorough taxonomy and a systematic review and highlights future research topics.
III-B2 Our Contributions
The main contributions of this paper are:
  • We conduct a systematic review that provided a thorough examination of edge AI research based on many applications, highlighting current trends and possible directions for the future.
@@ IV Review Methodology
This article does a systematic review to categorize studies that are pertinent to this study domain or discuss specific research questions on ”Edge AI”. In this article, we used the guidelines established by Kitchenham et al. [72, 73, 74] to provide a comprehensive review of Edge AI. The review is the optimal and reliable approach to documenting and analyse current research works. The systematic approach enables researchers to carefully analyse the positive and negative aspects of recent studies, conduct a thorough examination to identify potential gaps in research and future trends and difficulties, and provide a solid foundation and context for establishing a new study field. Furthermore, the complete research approach is presented in Fig. 5, which represents the structure of the process that was used in the systematic review research.
@@ IV-A Design and Plan of Review
The review procedure illustrates the methodologies applied to conduct a systematic review to minimize the potential for biased research. Therefore, possessing a pre-established process is crucial. In the absence of a systematic methodology, researchers’ predispositions can influence the process of selecting and analyzing studies. This may result in the omission of crucial inquiries essential for a thorough analysis and comprehension of the subject matter. The review process encompasses the research inquiries, exploration approach, criteria for selecting studies, procedures for assessing quality, and techniques for extracting and synthesizing data [75].
@@ IV-B Research Questions
Determining the research queries is crucial in the method of planning to develop a strong systematic review. The design of the research problem requires a thorough examination of existing literature studies. The primary aim of the present systematic review is to thoroughly examine and evaluate the various methods and strategies being employed for edge intelligence or edge AI. Furthermore, in order to emphasise the research findings and effectively showcase the useful consequences, the research questions that follow have been defined in Table II.
@@
RQ1 What are the fundamental techniques and strategies in Edge AI, and how do they impact data processing efficiency for different applications? This research question aims to explain the effect of various approaches on the efficacy of Edge AI in real-world applications. Methodologies and Strategies Sections 5.1, 5.2, and 5.4
RQ2 What strategies can be used to optimize heuristic and meta-heuristic algorithms for enhanced performance and resource management in Edge AI applications? The question evaluates the efficacy of heuristic and meta-heuristic approaches in improving performance and resource management inside Edge AI systems. Heuristic and Meta-Heuristic Methods Sections 5.4.1 and 5.4.2
RQ3 What are the challenges and solutions for using machine learning methodologies in Edge AI for real-time data processing? The research question emphasizes addressing the challenges encountered in the integration of machine learning techniques into Edge AI and possible solutions. Machine Learning Techniques Section 5.4.3
RQ4 How can edge AI applications, especially those that use dynamic environments, implement deep reinforcement learning to improve decision-making? This question examines the function of deep reinforcement learning in enhancing real-time decision-making inside Edge AI environments. Deep Reinforcement Learning Section 5.4.4
RQ5 How can Edge AI increase scalability and resource efficiency with less computational power and storage size? This question addresses the need for effective techniques for managing scalability and resources in Edge AI environments. Resource Management and Heterogeneity Sections 5.7 and 5.11.2
RQ6 What are the principal issues in resource allocation and task distribution for Edge AI applications, and how can they be mitigated? This question tries to determine challenges in the deployment of Edge AI and provide methods for efficient resource management. Resource Provisioning and Workload Distribution Sections 5.5.1 and 5.5.4
RQ7 What effects do architectural choices have on the scalability and performance of AI applications running on the edge? This question evaluates the efficacy of various architectural methodologies in Edge AI systems. Architectural Comparisons Section 6.2
RQ8 How does Edge AI resource management, including application placement and workload prediction, depend on important factors? This question explores essential components that enhance efficient resource management techniques in Edge AI. Resource Management Section 5.5 and 6.5
RQ9 What are the main challenges to federated learning and how might it enhance data sharing and communication among distributed Edge AI devices? Federated learning is potential for distributed Edge AI applications, but data synchronization is addressed in this question. Federated Learning, Data sharing Section 8.2
RQ10 What are the most important factors to consider when deciding between cloud, fog, and edge AI infrastructure, and how do these factors impact application performance? This question examines factors impacting infrastructure decisions in Edge AI and its effects on performance. Infrastructure Selection Section 5.1 and 6.1
RQ11 When it comes to real-time applications, how can edge AI provide data privacy and security, especially in vulnerable domains ? The significance and possible solutions for data privacy and security in Edge AI applications are the primary aims of this question. Security and Privacy Sections 5.4 and 5.5
@@ IV-C2 Search Query
A comprehensive search was conducted utilising Logical OR/AND operators to connect the keywords, concepts, synonyms, and abbreviations. The initial phase entails conducting an automated search using predetermined keywords that align with the study topics of this systematic literature review (SLR). The keywords used are [(((“Edge AI” OR “Edge Intelligence” OR “Edge Computing”) AND (“Machine Learning” OR “Deep Learning” OR “Reinforcement Learning”)) AND (“Resource Management” OR “Resource Allocation”)) OR (“Cloud Computing” OR “Fog Computing”) OR (“Application Placement” OR “Security” OR “Scheduling” OR “Simulation”)]. The search terms are obtained from the specified research topics and the framework of this systematic review in order to encompass the most significant and interrelated publications.
@@ V A Taxonomy of Edge AI
In this taxonomy, we identify and categorize the current studies for Edge AI. To create a new taxonomy, the first three authors carefully examine the contents of the 60 papers available and then obtain the research summary as explained in Section IV. We classify the obtained solutions under the necessary headings. Based on the current advancements and existing literature on edge AI, we have proposed a taxonomy, as shown in Fig. 7. Leveraging edge AI-driven edge computing systems can be optimized. The best example of this is the EdgeAI-based latency-reducing studies found in the literature. We generally examine literature studies under 11 subheadings: infrastructure, application architecture, IoT use cases, methods, resource management, ML model sizing, heterogeneity, security, scheduling, container migration, and container scaling.
@@ V-A Infrastructure
AI models and applications can be deployed in different infrastructures based on their target scenarios. The three most common infrastructures widely used are Cloud, Fog, and Edge.
@@ V-A1 Cloud
The essence of cloud computing is to pool resources with virtualization technology. Virtualization technology transforms a physical machine into several virtual machines, greatly changing the mode of application operation and deployment. The term ’cloud’ here refers to remote data centers [21]. Users usually connect to the servers in the data center through the network to use the computing resources here. The essence of cloud computing is to pool resources, allowing users to purchase resources according to their own requirements and greatly reducing resource waste.
@@ V-A2 Fog
As is well known, fog is closer to the ground than cloud, and the ground here refers to the user’s local device. Fog computing technology adopts distributed computing technology, which arranges computing resources on the side closer to the user than cloud computing [6]. It can be said that fog computing has broadened the network computing mode of cloud computing, widening computing capacity from the network center to the network edge, and thus more widely applied to numerous services. Fog computing is more widely distributed geographically and has greater mobility, making it suitable for an increasing number of intelligent devices that do not require extensive computation. For some time delay-sensitive applications such as real-time interaction systems, fog computing also has greater advantages.
Image, video and natural language processing (NLP) are the recently emerging fog computing applications. The placement and processing of images in fog computing is one of the most widely used fields of AI in research and industry. Its goal is to distinguish objects and people from each other and classify and distinguish photos based on image processing algorithms. Using fog computing in image processing-based applications can shorten response delay and improve service quality. In medical applications that require image processing accuracy and fast processing of medical data, deploying effective scheduling algorithms in foggy environments may be beneficial. According to other works, DL algorithms such as Generative Adversarial Networks (GAN) and Convolutional Neural Networks (CNN) can commonly be used in the field of image processing in fog.